### v2

Abstract
This paper presents GraphMend, a high-level compiler technique that eliminates FX graph breaks in PyTorch 2 programs. Although PyTorch 2 introduced TorchDynamo and TorchInductor to enable just-in-time graph compilation, unresolved dynamic control flow and unsupported Python constructs often fragment models into multiple FX graphs. These fragments force frequent fallbacks to eager mode, introduce costly CPU-to-GPU synchronizations, and reduce optimization opportunities. GraphMend addresses this limitation by analyzing and transforming source code before execution. Built on the Jaseci compilation framework, GraphMend introduces two code transformations that remove graph breaks due to dynamic control flow and Python side effects. This design allows PyTorch’s compilation pipeline to capture larger, uninterrupted FX graphs without requiring manual refactoring by developers. Evaluation across eight Hugging Face models shows that GraphMend removes graph breaks due to dynamic control flow and Python side effects, reducing the break count to 0 in 6 models and reducing it from 5 to 2 in another model. On NVIDIA RTX 3090 and A40 GPUs, GraphMend achieves up to 75% latency reductions and up to 8% higher end-to-end throughput. These results demonstrate that high-level code transformation is an effective complement to PyTorch’s dynamic JIT compilation pipeline, substantially improving both usability and performance.
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
PyTorch has emerged as one of the most widely adopted deep learning frameworks in both academia and industry, primarily due to its ease of use and flexibility Paszke et al. (2019). However, its general-purpose design incurs efficiency overheads, leaving PyTorch workloads less optimized than handcrafted GPU implementations. To bridge this gap, PyTorch 2 Ansel et al. (2024) introduced a Python Just-in-Time (JIT) compilation pipeline that translates model operations into optimized GPU code, achieving performance close to handcrafted implementations.
One of the central intermediate representations (IR) in this pipeline is the FX graph Reed et al. (2022). TorchDynamo intercepts and symbolically evaluates Python bytecode to create an FX graph. The FX graph is then passed to a backend compiler, by default TorchInductor Ansel et al. (2024), which applies a series of graph-level and kernel-level optimizations before lowering the computation to efficient kernels for GPU execution.
When tracing the FX graph, TorchDynamo often struggles with unsupported Python operations, which it handles by falling back to eager mode. These unsupported operations include dynamic control flow such as conditionals that depend on runtime values and certain Python side effects like I/O operations. In such cases, it inserts graph breaks, resulting in multiple disjoint FX graphs instead of a single unified one Zheng et al. (2023); Ansel et al. (2024).
These disjoint FX graphs run on the optimized backend, while the regions between them execute in standard PyTorch eager mode Paszke et al. (2019). Each fallback to eager mode forces GPU kernels to synchronize with the CPU to schedule subsequent operations, introducing significant overhead from repeated host-device communication. These synchronizations not only fragment execution but also reduce optimization opportunities, as each FX graph is compiled independently, preventing cross-graph fusion and global optimization Agrawal et al. (2019); Ansel et al. (2024).
Avoiding graph breaks is essential for realizing the full performance potential of PyTorch 2. Some of these breaks can be eliminated by transforming Python code into forms that use PyTorch-supported operations, allowing the compiler to capture more of the program as a single, continuous FX graph. However, the current compilation pipeline analyzes code at the bytecode level, operating instruction by instruction without understanding higher-level program structures such as control flow, data dependencies, and side effects. Our key insight is that analyzing and transforming higher-level program representations closer to the source code, and integrating this with bytecode-level analysis, enables a more holistic compilation strategy that exposes richer optimization opportunities.
1def f(x, y):
2 x_1 = x*2
3 y_1 = y*2
4 if x.sum() > 10: # <-- graph-break here
5 z = x_1 + y_1
6 else:
7 z = x_1 * y_1
8 return torch.relu(z)
Figure 1: A PyTorch forward pass with data-dependent control flow. The conditional branch on x.sum cannot be captured, causing a graph break. (See also Figure 2 for the transformed version that eliminates the break.)
Building on this insight, we introduce GraphMend, a high level compiler technique for PyTorch 2 that eliminates two major types of graph breaks, those arising from dynamic control flow and several Python side effects, through high-level program analysis and transformation. GraphMend extends the standard PyTorch 2 compilation pipeline with an AST-level transformation phase that restructures Python code before bytecode generation, improving traceability by eliminating graph breaks.
We implement GraphMend within the Jaseci framework, built on Jac language Mars et al. (2023); Mars (2025); Jaseci Labs (2025), which is a super-set of Python. The Jac compiler constructs an abstract syntax tree (AST) and a control flow graph (CFG), merging them into a unified intermediate representation (UniiR) Dantanarayana et al. (2025). Multiple compiler passes analyze and transform this IR, and GraphMend extends this pipeline with additional passes that detect and restructure code patterns prone to graph breaks. Specifically, it introduces two transformations, Predicated Dynamic Control Flow and Graph Epilogue Deferred Side Effects, which systematically eliminate these breaks.
We evaluate GraphMend on a benchmark suite of 8 Hugging Face Hugging Face (2025) models that trigger FX graph breaks. GraphMend removes all fixable breaks, reducing the break count to 0 in 6 models and from 5 to 2 in another, with 1 remaining unfixed. On NVIDIA RTX 3090 and A40 GPUs, it achieves 30–75% lower cold-start forward latency, 2.5–25% lower steady-state latency, and 5–8% higher end-to-end throughput.
Our contributions are, (1) We introduce a benchmark suite of eight Hugging Face models that naturally exhibit graph breaks, providing a realistic and reproducible testbed for evaluating PyTorch compilation. (2) We design and implement the GraphMend compilation technique, which automatically restructure high level code to eliminate common graph breaks. (3) We conduct comprehensive evaluations on this benchmark suite using real world models and hardware, demonstrating the effectiveness of our approach in reducing graph breaks and improving performance.
1def f(x, y):
2 x_1 = x*2
3 y_1 = y*2
4 cond = x.sum > 10
5 z_add = x_1 + y_1
6 z_mul = x_1 * y_1
7
8 # rewrite dynamic control flow using tensor select
9 z = torch.where(cond, z_add, z_mul)
10 return torch.relu(z)
Figure 2: Eliminating the graph break by rewriting data-dependent control flow with torch.where, keeping the computation in a single FX graph. Together with Figure 1, these illustrate how dataflow rewriting replaces dynamic Python branching with predicated tensor execution.
Figure 3: Profiled traces of forward pass execution across CPU and GPU. (a) Forward pass execution trace of code with graph breaks in Figure 1. (b) Forward pass execution trace of equivalent code with graph breaks fixed in Figure 2.
(a) Existing graph-break in Phi-4-mini instruct model built by Microsoft.
(b) Version of Phi-4-mini instruct model with the graph break fixed, without affecting program correctness.
Figure 4: (a) Existing graph-breaks in a real model and (b) how it can be fixed to produce a single contiguous FX graph.
2 Motivation
FX graph capturing:
TorchDynamo Ansel et al. (2024) captures graphs by symbolically interpreting Python bytecode, extracting PyTorch operations into FX graphs while falling back to eager mode for unsupported code.
FX graph breaks.
Let us consider the program in Figure 1, which has a conditional branch. When Dynamo tries to symbolically evaluate the function to capture FX graph, it cannot capture the conditional branch because the outcome depends on the runtime value of the tensor expression x.sum. Therefore, it inserts a graph break. As a result, the computation before the branch is recorded in one FX graph, the conditional is executed in eager mode, and the remaining operations are traced into a second FX graph Ansel et al. (2024); PyTorch Contributors (2025).
Impact of graph breaks.
When graph breaks occur, the function is divided into multiple FX graphs, each compiled separately into GPU kernels Ansel et al. (2024); Ghosh et al. (2025). Each FX graph forms a distinct torch.compiled region separated by segments of unsupported Python code PyTorch Contributors (2025).
During execution, GPU runs the first compiled region. Control then returns to the Python interpreter to execute the unsupported code in eager mode and switches back to GPU for the next compiled region Ansel et al. (2024). These context switches introduce CPU–GPU synchronization overhead and reduce GPU utilization.
Figure 3(a) illustrates this effect for the function in Figure 1. The trace shows two separate torch.compile regions divided by a device-to-host (D2H) memory transfer, where the GPU remains idle until the next CUDA graph launch. This fragmentation increases CPU scheduling overhead and prevents optimizations such as kernel fusion. When the break is removed and captured as a single FX graph, the execution trace in Figure 3(b) shows one continuous CUDA region with full GPU utilization and no intermediate D2H transfers. Kernel launches become streamlined, and synchronization overhead is eliminated.
Real world model example with graph breaks.
Phi-4-mini Instruct Microsoft (2025) is a lightweight open-source instruction-following model from Microsoft, available on Hugging Face as an efficient alternative to larger transformers. According to Hugging Face statistics Hugging Face (2025), the model recorded over 221k downloads in September 2025, reflecting strong community adoption.
To examine the impact of graph breaks in a widely used model, we profiled Phi-4-mini Instruct using PyTorch 2 and identified 5 graph breaks caused by data-dependent control flow. Figure 4(a) illustrates this case, where conditional logic depending on runtime tensor values introduces a graph break.
We rewrote the same control flow using PyTorch-compatible transformations with the torch.where operation to evaluate both branches within the FX graph while correctly handling side effects, as shown in Figure 4(b). After applying this modification, all graph breaks were eliminated. Profiling showed over a 50% latency reduction per forward pass in cold runs across multiple GPUs and more than a 5% improvement in steady-state runs compared to the version with graph breaks (further discussed in § 5), by reducing the impact of the graph breaks. This example demonstrates that targeted code transformations can effectively remove graph breaks and substantially improve performance.
3 Understanding and Eliminating Common Graph Breaks
1def fn(x):
2 x = torch.relu(x)
3 print("tensor:", x) # <-- graph-break here
4 return torch.sin(x)
Figure 5: Compiled function with a print statement. Dynamo inserts a graph break at the print statement. (See also Figure 6 for the transformed version that avoids the break.)
1def fn(x):
2 x = torch.relu(x)
3 to_print = "tensor:", x # <-- assign variable to print
4 y = torch.sin(x)
5 print(to_print) # <-- print at the end of tracing
6 return y
Figure 6: Fixing the graph break in Figure 5 by registering print as a reorderable logging function. The value to be printed is assigned to a buffer during tracing and flushed at the end.
Figure 7: GraphMend compiler integration in the Jac pipeline. The pipeline (top) lowers Python/Jac source code into a unified intermediate representation (UniiR) with AST, symbol table, and CFG. GraphMend (bottom) analyzes entry points, detects graph breaks, and applies AST transformations (Predicated Dynamic Control Flow and Graph-Epilogue Deferred Side Effects) to eliminate breaks before execution.
In this work, we focus on two common causes of graph breaks, outlining code transformations to eliminate them and explaining why bytecode-level fixes are challenging without higher-level IR context.
Data dependent operations.
Ansel et al. (2024); PyTorch Contributors (2025) In PyTorch 2, symbolic tracing builds FX graphs using FakeTensor objects, assuming control flow decisions can be inferred from static information such as tensor shapes or constants. When a condition depends on actual tensor data, (for example, if x.sum() > 0: in Figure 1), the compiler cannot safely evaluate it, so PyTorch inserts a graph break.
We can fix them by expressing data-dependent branches using tensor operations such as torch.where or torch.cond, which are supported on GPU. Figure 1-2 shows the rewritten version of the same function, where the use of torch.where eliminates the graph break.
Python Side-effects.
Side effecting operations such as print or logging calls also cause graph breaks because they interact with the Python runtime and cannot be represented in a functional graph. Deferring these effects until after the graph executes can remove the break while preserving program behavior. In Figure 5 and 6, we show how moving a print statement to the end of the function allows the entire function to remain in one FX graph.
Why automating these transformations at the bytecode level is challenging?
Automating these transformations must occur before PyTorch’s bytecode analysis, but performing safe transformations at the bytecode level is challenging since much of the high-level semantic information is lost during compilation.
When high-level constructs such as if/else or while loops are lowered to bytecode, they become low-level jump instructions like POP_JUMP_IF_FALSE and JUMP_BACKWARD. Since these opcodes are reused across different control-flow patterns, distinguishing loops from conditionals and identifying their bodies is non-trivial and often requires partial reverse engineering of the bytecode. Ensuring correctness during transformation further complicates this process.
Moreover, performing data-flow analysis at the bytecode level is difficult since Python uses a stack-based rather than a register-based model. While variable operations themselves are explicit, intermediate expression values are pushed and popped from the operand stack without explicit naming, obscuring the flow of computed values without symbolic stack simulation. This simulation adds complexity and overhead to the analysis, whereas analyses on higher-level AST representations with explicit variable bindings and control flow are far simpler.
Handling side effects at the bytecode level is also challenging. While delaying operations like built-in print or logger calls may be safe, Python’s dynamic nature allows users to redefine these functions. Determining whether a call refers to the original built-in or a user-defined override requires tracking symbol bindings, which is only possible with higher-level semantic context.
A higher-level IR that retains program structure and data dependencies is therefore better suited for automatically applying these transformations.
4 GraphMend Compiler Technique
In this section we introduce our compiler technique GraphMend, which automates these code transformations using AST and CFG level analyses and AST-transformations, to eliminate graph breaks. The implementation of this technique is shown in Figure 7 which is elaborated in the rest of this section.
4.1 Jac Compilation Pipeline
To implement our compiler technique, we build on the Jaseci framework Mars et al. (2023); Mars (2025); Dantanarayana et al. (2025), a compiler stack designed for advanced analysis and transformation of Python programs. Jaseci compiler accepts both Jac and Python code, parses them into an AST, and progressively lowers the representation into UniiR. As shown in the top half of Figure 7, UniiR unifies the AST, CFG and symbol table into a single structure, combining syntactic, semantic, and control-flow information.
Input: UniiR; TorchAttrTable (torch attributes $\mapsto$ {dynamic, static})
Output: Tags: GraphBreak[DynCtrlFl], GraphBreak[logger/print]
1
2foreach Dynamo Entry Point $r$ in UniiR do
3   Initialize worklist with successors of $r$
4 while worklist not empty do
5      Pop node $n$ from worklist; push unseen successors of $n$
6
    // 1) Dynamic control-flow via if-conditions
7    if $n$ is an IfStmt then
8       $E\leftarrow$ subexpressions of $\texttt{cond}(n)$
9       $dyn\leftarrow\textbf{false}$
10       foreach $e\in E$ do
11          if $e$ has a torch attribute then
             $a\leftarrow$ attribute of $e$
             // e.g., .sum
12             if TorchAttrTable[$a$] = dynamic then
13                $dyn\leftarrow\textbf{true}$; break
14
15
16
17       if not dyndyn then
18          if $E$ is input/tensor-value dependent (via SymbolTable + CFG) then
19             $dyn\leftarrow\textbf{true}$
20
21
22       if dyndyn then
23            Tag $n$ as GraphBreak[DynCtrlFl]
24
25
26
    // 2) Logger / print side effects
27    if $n$ is a call to print or logger.* then
28         Tag $n$ as GraphBreak[logger/print]
29
30
31
Algorithm 1 Fixable Graph-Break Detection
4.2 GraphMend Implementation
Our compiler technique consists of three main passes, illustrated in the bottom half of Figure 7.
Dynamo Entry Point Analysis Pass
To identify potential TorchDynamo compilation entry points, this pass traverses the AST and analyzes each node. A node is tagged when a function or an nn.Module object wrapped or decorated with torch.compile is encountered. If no occurrences of torch.compile are found in the codebase, GraphMend compilation is skipped. Otherwise, the tagged nodes serve as entry points for subsequent passes.
Graph Break Type Analysis Pass.
We implement a compound analysis pass on UniiR to identify potential graph-break locations, following the procedure in Algorithm 1. Starting from each Dynamo entry point, the pass traverses the CFG to locate program points that may induce graph breaks. For conditional branches, the pass inspects the subexpressions of the condition. If any involve Torch attributes known to be dynamic (e.g., .sum()), the branch is marked as dynamic control flow. Otherwise, a use-def analysis over the symbol table and CFG determines whether the condition depends on model inputs or tensor values; if so, it is also tagged as dynamic. In addition, side-effecting operations such as print and logger.* calls are detected and annotated as logging-induced graph breaks as in Algorithm 1. These annotations guide the subsequent transformation passes that eliminate graph breaks.
AST Transformation Pass
After identifying graph break locations, we modify the AST according to the graph break tags detected in the previous pass, after verifying the transformation is safe. Based on these tags, one of the following transformations is applied.
Predicated Dynamic Control Flow Transformation
Before performing code transformation to address data-dependent control flow that causes graph breaks, we first analyze each branch body to identify which variables are modified and which functions are invoked in the if_true and if_false branches. If both branches produce the same output variables, their results can first be computed separately using two different temporary variables, and then combined with a torch.where expression that selects values based on the condition. If both branches call the same function, the function is moved outside the branches and executed only once, after its input arguments are merged using torch.where.
The transformation proceeds by rewriting the AST to handle data-dependent control-flow graph breaks. The branch condition is first evaluated and stored as a predicate (pred). Each branch is then expressed as a dataflow computation producing temporary intermediate results, which are merged using torch.where(pred, e1, e2). After this reconciliation step, we handle common function calls ensuring that both their inputs and side-effect-free semantics are preserved. Finally, the original if and else nodes are removed, and their child nodes are reconnected to the parent, yielding a transformed AST in which both execution paths are represented explicitly as dataflow expressions. This process, illustrated in Figure 7, eliminates data-dependent control-flow constructs and enables predicated, differentiable execution compatible with graph compilation.
Table 1: Models in benchmark suite
Benchmark Model Name Graph break count Graph break reasons Total Operators
biogpt Luo et al. (2022) 2 logger calls 408
blenderbot-400M-distill Meta AI (2020) 3 logger calls 489
flan-t5-large Chung et al. (2022) 3 logger calls 2130
longformer-base-409 Beltagy et al. (2020) 5 logger calls, tensor.item() call 913
moe-minicpm-x4-base babybirdprd (2025) 15 Dynamic shape operator 280
Phi-4-mini-instruct Microsoft (2025) 5 Dynamic control flow 1541
Qwen-Audio-Chat Chu et al. (2023) 2 Dynamic control flow 1647
tiny-random-PegasusForCausalLM Face (2025) 2 logger calls 64
Deferred Side Effect Transformation
Before applying this transformation, we perform analysis to ensure that print and logger statements correspond to the built-in implementations rather than user-overloaded versions. This prevents interference with user-defined behavior or unintended semantic changes. Once verified, side-effecting calls such as print or logger are rewritten into deferred buffer assignments. Specifically, the value to be printed/ logged is stored in a dedicated runtime buffer together with a predicate flag indicating whether the branch containing the statement was executed. After the FX graph finishes execution, the buffer is flushed: each stored message is printed or logged only if its flag is true. This keeps side-effecting operations in the right order and with the same behavior, but safely outside the traced region.
At the end of the Jac compiler pipeline, Python bytecode is generated for interpretation, enabling TorchDynamo to seamlessly operate on the transformed code.
5 Evaluation
Table 2: Graph break counts in the original model and fix rates achieved by applying GraphMend across the benchmark suite.
Benchmark Model Graph Breaks Fixed (%)
biogpt 2 100
blenderbot-400M-distill 3 100
flan-t5-large 3 100
longformer-base-4096 5 40
moe-minicpm-x4-base 15 0
Phi-4-mini-instruct 5 100
Qwen-Audio-Chat 2 100
tiny-random-PegasusForCausalLM 2 100
Figure 8: Latency improvements from GraphMend (a) Cold-start latency reductions (b) Steady-state latency reductions across benchmark models on RTX 3090 and A40 GPUs.
Figure 9: Relative throughput improvement across benchmark suite
(a) Profiler tracer for one iteration of warm runs for Qwen-Audio-Chat original model before fixing the graph break run in A40 GPU
(b) Profiler tracer for one iteration of warm runs for Qwen-Audio-Chat model after fixing the graph break run in A40 GPU
Figure 10: (a) Profile traces.
(a) Steady State - Original Model
(b) Cold Start - Original Model
(c) Steady State - Fixed Model
(d) Cold Start - Fixed Model
Figure 11: CPU/GPU activity time traces for Qwen-Audio-Chat model run on A40 GPU.
We evaluate GraphMend on real-world models that exhibit graph breaks. Our goals are to quantify (i) how often GraphMend removes fixable breaks and (ii) the end-to-end improvement after those fixes.
Experimental Setup
To implement our compiler technique, we build on Jaseci framework Mars et al. (2023); Mars (2025). Experiments are conducted on NVIDIA RTX 3090 and A40 GPUs using TorchInductor as the backend compiler.
Benchmark Suite
To construct our benchmark suite, we (1) randomly sample 65 Hugging Face Hugging Face (2025) models spanning diverse architectures and task categories; (2) test each model to identify whether it triggers FX graph breaks; and (3) retain all 8 models that exhibit graph breaks for evaluation. Table 1 summarizes the resulting suite, including the number and causes of graph breaks and the total operator count. A similar methodology was used in prior work Ansel et al. (2024).
Profiling Methodology
We profile the forward pass using PyTorch Profiler pyt (2021). We evaluate the original and graph-break–fixed versions of each model, running seven iterations with one cold start and six warm runs. The profiler records CPU, CUDA, and operator-level events, which are then analyzed using Chrome’s trace viewer.
5.1 GraphMend’s Ability to Fix Graph Breaks
First, we evaluate GraphMend’s effectiveness in fixing graph breaks across the benchmark suite. As shown in Table 2, GraphMend fully eliminates graph breaks in 6 models and resolves some graph breaks in 2 others. Graph breaks introduced by tensor.item() calls and dynamic-shape operators remain unfixed, as they cannot be resolved through our code transformations. For the performance analysis that follows, we report results on the 7 models where graph breaks were fully or partially eliminated, excluding moe-minicpm-x4-base.
5.2 Latency Improvement per Forward Pass Using GraphMend to Fix Graph Breaks
We measure the relative latency improvement as the percentage reduction in per-forward-pass latency after applying GraphMend fixes compared to the version with graph breaks. Results are reported for the models using PyTorch Profiler data from both cold and warm runs.
Cold Start Latency Improvement
Figure 8(a) shows results on RTX 3090 and A40 GPUs. Across models, we observe a 30–75% reduction in latency during cold runs. At cold start, each graph break introduces significant overhead as separate CUDA graphs must be recorded and cached. Eliminating these breaks reduces capture operations to a single (or fewer) CUDA graph, minimizing overhead and yielding larger per forward pass improvements. A detailed analysis appears in  § 5.4.
Steady-State Latency Improvement
Figure 8(b) shows steady state results on RTX 3090 and A40 GPUs, with 2.5–25% latency reductions across models. After the cold run, cached CUDA graphs eliminate capture overhead, but remaining graph breaks still cause GPU to CPU switches, adding latency. Removing these breaks with GraphMend eliminates such switches, improving forward pass performance.
Notably, tiny-random-PegasusForCausalLM shows over 25% improvement. Since this model is very small (64 operators), the fixed costs from graph breaks (e.g. host–device synchronization, and extra graph launches) take up a large portion of the runtime. Removing these breaks therefore gives a huge speedup per forward pass.
5.3 Relative Throughput Improvement by Fixing Graph Breaks
We measure the relative throughput improvement as the percentage increase in output tokens per second after applying GraphMend fixes compared to the version with graph breaks.
We measure relative throughput for all models on RTX 3090 and A40 GPUs. Figure 8(b) shows 5–8% improvements, consistent with the latency reductions in Section 5.2. As each forward pass becomes faster, overall tokens per second increase. The variation of improvement across models reflects how much a graph break affects the forward path: when the impact of a break is small, the corresponding throughput gain is also limited.
For example, in both Phi-4-mini-instruct and Qwen-Audio-Chat, the root cause of the breaks is dynamic control flow. We observe a 7.5%–8% improvement for Qwen-Audio-Chat, compared to 5%–6% for Phi-4-mini-instruct. The difference arises because the break in Phi-4-mini-instruct occurs early in the forward function, while in Qwen-Audio-Chat it occurs mid-function, where its impact is greater. Hence, a break’s location determines its effect, fixing some yields larger throughput gains than others.
tiny-random-PegasusForCausalLM achieves more than 25% lower forward-pass latency, but throughput gains are smaller. This is because throughput also depends on prefill, decode, CPU tasks (e.g., sampling, tokenization), and data transfers, which are unaffected. For such a small model, these fixed costs dominate, so removing the break speeds up GPU compute but has less impact on end-to-end tokens/s, especially when decode dominates generation.
5.4 Overhead Analysis of Graph Breaks
We analyze how fixing graph breaks with GraphMend reduces overhead using the Qwen Audio Chat model on an A40 GPU, which contains two graph breaks.
For a consistent comparison, all activity plots (Figures 11(a)11(d)) show the time interval between the first and second CUDA graph executions in the original model; the same interval is used for the fixed model.
Steady-state run overhead analysis
Figure 10(a) shows the profiler trace for the original model. Execution is split into 3 CUDA graphs due to the 2 graph breaks, with idle GPU periods between the first and second FX graph executions. This is clearly visible in Figure 11(a), GPU becomes idle while CPU handles control flow in eager mode, including a D2H memcpy triggered by dynamic control flow. This CPU fallback introduces an overhead.
After applying GraphMend replacing torch.where, the resulting trace (Figure 10(b)) shows one continuous CUDA graph. As confirmed in Figure 11(c), GPU activity remains uninterrupted and CPU fallbacks disappear. Thus, GraphMend eliminates the overhead of CPU–GPU context switching during steady-state runs.
Cold run overhead analysis
Cold runs introduce even larger overheads. In the original model (Figure 11(b)), each graph break triggers a separate CUDA Graph recording and caching, causing long GPU idle phases and heavy CPU activity. GPU idle time dominates as the CPU handles repeated captures. In contrast, the fixed model (Figure 11(d)) performs a single capture, minimizing GPU idle time and CPU work. Consequently, GraphMend yields larger overhead reductions in cold runs, explaining the substantial cold-start latency gains per forward pass shown in Figure 8(a).
5.5 GPU Runtime Kernel Statistics Analysis
We examine how fixing graph breaks affects GPU kernel execution and optimization. On Phi-4-mini-instruct model (A40 GPU), the profiler reports 404 kernels for the original model and 393 for the fixed one; a reduction reflecting increased fusion and fewer launches as larger regions are captured.
As shown in Figure 12, three small kernels in the original model merge into a single fused kernel after fixing. Kernel reordering also appears, as the unified CUDA Graph enables better packing and scheduling of independent regions, reducing stalls and improving locality.
Graph breaks previously forced separate captures, preventing cross-break fusion and limiting scheduling windows. Their removal allows broader fusion, redundant launch elimination, and more flexible reordering; reducing kernel gaps and boosting device utilization.
In summary, fixing graph breaks with GraphMend not only reduces launch overhead but also exposes new optimization opportunities, underscoring the importance of eliminating such breaks.
Figure 12: Kernel Fusion and Reordering Visualization for Phi-4-mini-instruct model.
6 Related Work
Table 3: Comparison of GraphMend with related transformation frameworks.
Feature AutoGraph MagPy TensorSSA GraphMend
Moldovan et al. (2019) Zhang et al. (2024) Ma et al. (2024)
[…]
Fully compatible with PyTorch 2 pipeline $cellcolorlightred \cellcolorlightred \cellcolorlightgreen\checkmark$
Handling side effects compatible with PyTorch $cellcolorlightred \cellcolorlightred \cellcolorlightgreen\checkmark$
[…]
PyTorch has long sought to balance performance with Python’s flexibility. TorchScript’s tracing and scripting enabled ahead-of-time graph capture but struggled with dynamic control flow and unsupported constructs PyTorch-TorchScript Team (2025); PyTorch Team (2025). PyTorch 2 introduced TorchDynamo for bytecode-level graph capture Ansel et al. (2024); Paszke et al. (2019), yet data-dependent branches and Python side effects still cause frequent graph breaks. GraphMend mitigates these by transforming the Python AST before execution.
While AutoGraph Moldovan et al. (2019) generates computation graphs in TensorFlow through source-level transformations, its approach does not generalize to PyTorch 2, which captures graphs at the bytecode level. In contrast, GraphMend performs lightweight AST-level transformations integrated directly into the PyTorch 2 compilation pipeline, improving symbolic evaluation and graph completeness. MagPy Zhang et al. (2024) takes a different route by replacing TorchDynamo and reconstructing graphs from runtime traces, whereas GraphMend preserves the standard compilation flow with minimal overhead. TensorSSA Ma et al. (2024) introduces a custom SSA-based IR to manage tensor mutations and control flow but adds significant infrastructure and omits Python side effects handling. GraphMend instead applies predicate-based SSA-inspired rewrites directly on the Python AST, avoiding extra IR layers while staying fully compatible with PyTorch’s ecosystem.
More broadly, GraphMend’s source-level rewriting follows the tradition of semantics-preserving transformations in program optimization, as exemplified by Halide Ragan-Kelley et al. (2013), Lantern Wang et al. (2018), and DLVM Zhao et al. (2017). Like these systems, GraphMend leverages structural code analysis for performance gains but is uniquely tailored to PyTorch, enabling TorchDynamo to capture larger, uninterrupted graphs. General compiler infrastructures such as XLA Community (2023), TVM Chen et al. (2018), and MLIR Lattner et al. (2020) provide IR-based optimization, while Glow Rotem et al. (2018), TASO Jia et al. (2019), and ONNX Runtime Microsoft (2018) extend portability across hardware and frameworks. These systems assume large graphs are already available, whereas GraphMend restructures PyTorch code to maximize graph continuity so their optimizations can be applied more effectively.
7 Conclusion
We present GraphMend, a source-level compiler technique for PyTorch that systematically eliminates common FX graph breaks before execution. By restructuring code constructs such as dynamic control flow and Python side effects (e.g., I/O operations), GraphMend reduces fragmentation in the compilation pipeline and enables TorchDynamo to capture larger, unified graphs. Our evaluation across a diverse suite of Hugging Face models demonstrates that GraphMend removes nearly all fixable graph breaks and consistently improves runtime efficiency, achieving substantial reductions in latency and increases in throughput on modern GPUs.